OpenStudy (anonymous):

Taylor is making a cookie cake that has a diameter of 13 in. What is the area of a fourth of the cookie cake? A)36.97 B)36.25 C)33.18 D)38.4

OpenStudy (anonymous):

total area / 4 total area = pi x r^2 = pi x 6.5^2 = 132.73229 total area / 4 132 / 4 = 33.1830725 your answer is C

OpenStudy (anonymous):

Thank You (: